Find the total surface area, in square inches, of a cylinder with a radius r = 8 in. and height h = 5 in.
1
Expert's answer
2011-12-09T09:11:00-0500
The total surface area consists of cilinder basis area

S1 = 2*pi*R^2

and cilinder lateral surface area

S2 = 2*pi*R*H

Hence,
S = S1 + S2 = 2*pi*R^2 + 2*pi*R*H = 2*pi*8^2 + 2*pi*8*5 ≈ 653.4513 square inches.